# Extended Resolution Clause Learning via Dual Implication Points

### v2

@@ 5.1. An example with a grid Tseitin principle
Given a graph where every vertex has a charge, a number which is 00 or $1$, a Tseitin formula [Tse83] is created by considering one variable per edge, and adding one constraint per vertex $v$ expressing that the sum of the variables of all edges incident to $v$ modulo 2 is equal to its charge. The CNF version that we consider converts each xor constraint into clauses by simple enumeration of falsifying assignments. It is easy to see that the formula is unsatisfiable if and only if the sum of all charges is even. Here we consider as a graph the $3\times 3$-grid depicted in Figure 4 and only one vertex has charge 1, so the clauses are unsatisfiable.
@@
$e_{1}\oplus e_{3}$$=$ 1 $e_{5}\oplus e_{7}\oplus e_{10}$$=$ 0
$e_{1}\oplus e_{2}\oplus e_{4}$$=$ 0 $e_{8}\oplus e_{11}$$=$ 0
$e_{2}\oplus e_{5}$$=$ 0 $e_{9}\oplus e_{11}\oplus e_{12}$$=$ 0
$e_{3}\oplus e_{6}\oplus e_{8}$$=$ 0 $e_{10}\oplus e_{12}$$=$ 0
$e_{4}\oplus e_{6}\oplus e_{7}\oplus e_{9}$$=$ 0
@@
$e_{1}$evel 1:$\overline{e_{3}}$$e_{2}$evel 2:$\overline{e_{4}}$$e_{5}$$e_{6}$evel 3:$e_{8}$$e_{11}$$e_{7}$evel 4:$\overline{e_{9}}$$e_{12}$$\overline{e_{10}}$$\perp$(a):
$e_{1}$evel 1:$\overline{e_{3}}$$e_{2}$evel 2:$\overline{e_{4}}$$e_{5}$$e_{6}$evel 3:$e_{8}$$e_{11}$$\overline{x}$$\overline{e_{7}}$$e_{10}$$e_{9}$$\overline{e_{12}}$$\perp$(b):
@@ 5.2. Extending CDCL with Dual Implication Points
It would make no sense to backtrack to level $\ell_{C}$ when $\ell_{C}<\ell_{D}$ since then neither $\lnot z$ nor $\lnot f$ would be propagated. However, another possible strategy would be to backtrack to the maximum of the decision levels $\ell_{C}$ and $\ell_{D}$. This would mean $\lnot z$ and $\lnot f$ are both propagated. The disadvantage of this when $\ell_{C}>\ell_{D}$ is that it would mean $\lnot z$ is asserted by the post-DIP clause at level $\ell_{C}$, whereas it could have been propagated at the previous decision level $\ell_{D}$. This breaks a usual invariant for CDCL solvers. This would only be possible in a solver that permits chronological backtracking [NR18, MB19]; xMapleLCM, however, does not support this.
@@ 6. Experimental Evaluation
We have implemented the DIP-based clause learning schemes described in Section 5.2 on top of the xMapleLCM ER framework [^3] All sources used for this evaluation can be found in https://github.com/chjon/xMapleSAT/tree/main.. We started our experimental evaluation running a variant of DIP-based learning on benchmarks of the SAT Competition [BHI+23] and comparing it to MapleLCM [LLX+17], the CDCL SAT solver on which it is based. Even though there did not seem to be a systematic improvement on all benchmarks, a few families with important speedups were identified. We start this section by analyzing the impact of DIP-based learning on these families and then move to final considerations about the performance on the overall SAT competition benchmarks.
@@
After that, we report on the performance of a variety of state-of-the-art solvers, each with some distinguished characteristic: Kissat 4.0.2 [BFFH20] (an extremely efficient CDCL solver), CryptoMiniSat 5.12.1 [SNC09] (support for XOR reasoning), SBVA-CaDiCaL [HGH23] (introduction of new variables via SBVA and winner of the main track of the 2023 SAT Competition), GlucosER [AKS10] (extended-resolution based CDCL solver) and the best available configuration for xMapleLCM.
@@ 6.1. Tseitin Formulas
These formulas have already been described in Section 5.1. In this section, we consider three types of unsatisfiable formulas generated by CNFgen [LENV17]: (i) Tseitin formulas on the grid, where the graph is a rectangular grid and each vertex is connected to its 4 neighbors, except for the vertices on the boundaries of the grid which have fewer neighbors, (ii) 4-regular Tseitin formulas, where the graph is a $4$-regular random one, and (iii) 6-regular Tseitin formulas built on top of $6$-regular random graphs. The time limit was set to 600 seconds per instance.
@@
In order to understand whether the behavior of xMapleLCM was polynomial for these formulas, we generated more challenging Tseitin formulas on the grid. Results can be seen in the left plot of Figure 7, with logarithmic scale on the $y$ axis. There is little doubt that the runtime of our solver ends up being exponential w.r.t. the size of the problem. However, we went one step further and studied how large the generated DRAT proofs were. On the right plot, we show the number of resolution steps in the proof. Despite having no theoretical support for that, it seems that although the solver takes exponential time in finding a proof, its size might be polynomial w.r.t. the problem size. That would indicate that our problem is using search heuristics that are not good enough to quickly find a short proof.
@@ 6.2. Xorified Random k-XOR Formulas
Another family where DIP-based systems perform very well are random $k$-xor formulas, where xorification [BS09] has been applied, i.e., replacing variables by xors of fresh variables. We used CNFgen to obtain 2 sets of formulas: xor constraints of length 3 applying xorification with 2 and 3 variables. All formulas we consider are unsatisfiable, the number of variables before xorification is equal to the number of clauses, and we increase this number to get progressively more difficult formulas. A time limit of 600 seconds was used.
@@ 6.3. Matching of Properly Intersecting Intervals
We are given a sequence of numbers $(a_{1},\ldots,a_{n})$, initially all set to zero, and a set of operations, each consisting of assigning a certain number to a contiguous subsequence of the $a_{i}$’s, defined by an interval. The goal is to perform each operation exactly once while maximizing the number of pairs of consecutive numbers $(a_{i},a_{i+1})$ that are different at the end of the process. Given some additional conditions, this can be formulated as maximum bipartite matching problem on a certain graph. By removing some edges from this graph, an unsatisfiable problem is generated. A more detailed description can be found in [BHI+23].
We downloaded the 23 instances submitted to the SAT Competition, using the Global Benchmark Database [IS19], and observed that no system without extended-resolution reasoning could solve any benchmark in a time limit of 2 hours. We used a much larger timeout than with the two previous families because for this family we are not able to construct increasingly larger benchmarks and see how the solvers scale.
Table 1 reports on our experimental evaluation. In the first block we can see that the random and heuristic DIP choices do significantly improve upon our baseline solver. The second block row shows how the behavior of the heuristic configuration can be further improved by increasing the threshold of minimum DIP occurrences to 50. On the contrary, a smaller threshold of 5 significantly degrades the performance, whereas only learning the post-DIP clause slightly slows it down. Finally, the third block compares our best xMapleLCM configuration with state-of-the-art solvers. It is remarkable that, apart from GlucosER, no system is able to process any instance within the given time limit. Hence, the use of extended resolution is crucial in this set of benchmarks. Our comparison shows that GlucosER and xMapleLCM solve the same number of instances, the former being only slightly faster.
@@ 6.4. Performance Analysis on SAT Competition Benchmarks
In the left cactus plot of Figure 9 we can see that our baseline DIP-based xMapleLCM solver outperforms the CDCL-based MapleLCM that it is based on. When we analyzed the concrete benchmarks where our DIP-based solver outperforms MapleLCM and other CDCL-based solvers like Kissat, we observed that for all of them the percentage of decisions on extended variables is surprisingly large (over 10% of the decisions), whereas for the remaining benchmarks it is very low (in 70% of the benchmarks less than 1% of decisions are on extended variables). Moreover, this happens no matter how many initial variables there are, even though in problems with few number of initial variables the introduction of a few extended variables could quickly dominate the decision heuristic.
This is remarkable since it shows that already existing decision heuristics like VSIDS or LRB somehow infer whether the newly introduced variables improve the behavior of the system. This is why we implemented, on top of our baseline DIP system, a procedure that computes the percentage of decisions on extended variables. If after a given number of conflicts it is still lower than 3%, it discontinues DIP learning and performs 1-UIP learning from that moment onwards. This is a very preliminary step in the direction of trying to automate the decision of whether to apply DIP reasoning or not, but the outcome, which is found on the left cactus plot of Figure 9, is very positive, showing that a solver disabling DIP reasoning (xMapleLCM-disabling in the plot) when the percentage of decisions on extended variables is low is still able to solve Tseitin, random $k$-xor and Interval benchmarks, but the overall performance is not significantly worsened on the remaining instances. Finally, the right cactus plot shows the impact of changing the type of DIP into the disabling solver: the heuristic and middle DIPs, the latter being used in the baseline configuration, are much better than random and closest ones.
@@
  • We have identified three families of benchmarks for which the dynamic addition of extension variables improves upon CDCL in a systematic manner. This is an improvement w.r.t. the results in the GlucosER paper [AKS10], where a few initial benchmarks were identified but it was not proved that extended-resolution methods always outperform CDCL on increasingly more complex instances.
  • We have proved that, even though the way extended variables are identified is completely different in xMapleLCM and GlucosER, they both behave well on the same instances. This is remarkable because our preliminary experiments revealed that randomly choosing arbitrary pairs of (non-DIP) literals to create definitions does not provide any speedup w.r.t. CDCL. Hence, both methods are able to identify semantic properties about concrete literals that make them suitable for creating an extended variable definition.
@@ 7. Conclusions and Future Work
We introduce a novel extended resolution clause learning (ERCL) algorithm that, when implemented on top of a CDCL solver, turns out to be beneficial for a variety of problems, in particular Tseitin, random $k$-xor and interval matching formulas. This is remarkable since automating powerful proof systems is well-known to be a difficult task. Somewhat surprisingly, the only previously existing attempt to incorporate ER into CDCL performs reasonably well on the same instances. We have hence indentified classes of formulas for which extended resolution can be automated in at least two different ways and still outperform CDCL. Considering the different nature of the two methods, this clearly deserves further study. Further, we also introduce a new heuristic that allows our ERCL solver xMapleSAT to perform similarly to the baseline CDCL solver MapleLCM, thus being able to get the best of both worlds, i.e., the benefit of ERCL, without sacrificing performance of the CDCL solver on, say, SAT Competition Main Track 2023 instances.
As future work, we plan to investigate a variety of machine learning based heuristics (e.g., branching) specialized for the ERCL method. Also, the use of $k$-IPs with $k>2$ is part of our next steps. On the theoretical side, challenging questions like determining whether DIP- or $k$-IP based ER simulates unrestricted ER are going to be central to our research efforts.

### v3

@@ 5.1. An example with a grid Tseitin principle
Given a graph where every vertex has a charge, a number which is 0 or $1$, a Tseitin formula [tseitin1983complexity] is created by considering one variable per edge, and adding one constraint per vertex $v$ expressing that the sum of the variables of all edges incident to $v$ modulo 2 is equal to its charge. The CNF version that we consider converts each xor constraint into clauses by simple enumeration of falsifying assignments. It is easy to see that the formula is unsatisfiable if and only if the sum of all charges is odd. Here we consider as a graph the $3\times 3$-grid depicted in Figure 5 and only one vertex has charge 1, so the clauses are unsatisfiable.
@@
$e_{1}\oplus e_{3}$== 1 $e_{5}\oplus e_{7}\oplus e_{10}$== 0
$e_{1}\oplus e_{2}\oplus e_{4}$== 0 $e_{8}\oplus e_{11}$== 0
$e_{2}\oplus e_{5}$== 0 $e_{9}\oplus e_{11}\oplus e_{12}$== 0
$e_{3}\oplus e_{6}\oplus e_{8}$== 0 $e_{10}\oplus e_{12}$== 0
$e_{4}\oplus e_{6}\oplus e_{7}\oplus e_{9}$== 0
@@
[…]
@@ 5.2. Extending CDCL with Dual Implication Points
It would make no sense to backtrack to level $\ell_{C}$ when $\ell_{C}<\ell_{D}$ since then neither $\lnot z$ nor $\lnot f$ would be propagated. However, another possible strategy would be to backtrack to the maximum of the decision levels $\ell_{C}$ and $\ell_{D}$. This would mean $\lnot z$ and $\lnot f$ are both propagated. The disadvantage of this when $\ell_{C}>\ell_{D}$ is that it would mean $\lnot z$ is asserted by the post-DIP clause at level $\ell_{C}$, whereas it could have been propagated at the previous decision level $\ell_{D}$. This breaks a usual invariant for CDCL solvers. This would only be possible in a solver that permits chronological backtracking [NadelRychin:chronological, MohleBiere:backtracking]; xMapleLCM, however, does not support this.
@@ 6. Experimental Evaluation
We have implemented the DIP-based clause learning schemes described in Section 5.2 on top of the xMapleLCM ER framework [^4] All sources used for this evaluation can be found in https://github.com/chjon/xMapleSAT/tree/main.. We started our experimental evaluation running a variant of DIP-based learning on benchmarks of the SAT Competition [SATcomp2023] and comparing it to MapleLCM [MapleLCM], the CDCL SAT solver on which it is based. Even though there did not seem to be a systematic improvement on all benchmarks, a few families with important speedups were identified. We start this section by analyzing the impact of DIP-based learning on these families and then move to final considerations about the performance on the overall SAT competition benchmarks.
@@
After that, we report on the performance of a variety of state-of-the-art solvers, each with some distinguished characteristic: Kissat 4.0.2 [kissat] (an extremely efficient CDCL solver), CryptoMiniSat 5.12.1 [cryptominisat] (support for XOR reasoning), SBVA-CaDiCaL [SBVA] (introduction of new variables via SBVA and winner of the main track of the 2023 SAT Competition), GlucosER [Audemard2010ARO] (extended-resolution based CDCL solver) and the best available configuration for xMapleLCM.
@@ 6.1. Tseitin Formulas
These formulas have already been described in Section 5.1. In this section, we consider three types of unsatisfiable formulas generated by CNFgen [CNFGen]: (i) Tseitin formulas on the grid, where the graph is a rectangular grid and each vertex is connected to its 4 neighbors, except for the vertices on the boundaries of the grid which have fewer neighbors, (ii) 4-regular Tseitin formulas, where the graph is a $4$-regular random one, and (iii) 6-regular Tseitin formulas built on top of $6$-regular random graphs. The time limit was set to 600 seconds per instance.
@@
In order to understand whether the behavior of xMapleLCM was polynomial for these formulas, we generated more challenging Tseitin formulas on the grid. Results can be seen in the left plot of Figure 8, with logarithmic scale on the $y$ axis. There is little doubt that the runtime of our solver ends up being exponential w.r.t. the size of the problem. However, we went one step further and studied how large the generated trimmed DRAT proofs were. On the right plot, we show the number of resolution steps in the proof. Despite having no theoretical support for that, it seems that although the solver takes exponential time in finding a proof, its size might be polynomial w.r.t. the problem size. That would indicate that our problem is using search heuristics that are not good enough to quickly find a short proof.
@@ 6.2. Xorified Random k-XOR Formulas
Another family where DIP-based systems perform very well are random $k$-xor formulas, where xorification [BenSasson:sizespace] has been applied, i.e., replacing variables by xors of fresh variables. We used CNFgen to obtain 2 sets of formulas: xor constraints of length 3 applying xorification with 2 and 3 variables. All formulas we consider are unsatisfiable, the number of variables before xorification is equal to the number of clauses, and we increase this number to get progressively more difficult formulas. A time limit of 600 seconds was used.
@@ 6.3. Matching of Properly Intersecting Intervals
We are given a sequence of numbers $(a_{1},\ldots,a_{n})$, initially all set to zero, and a set of operations, each consisting of assigning a certain number to a contiguous subsequence of the $a_{i}$’s, defined by an interval. The goal is to perform each operation exactly once while maximizing the number of pairs of consecutive numbers $(a_{i},a_{i+1})$ that are different at the end of the process. Given some additional conditions, this can be formulated as maximum bipartite matching problem on a certain graph. By removing some edges from this graph, an unsatisfiable problem is generated. A more detailed description can be found in [SATcomp2023].
We downloaded the 23 instances submitted to the SAT Competition, using the Global Benchmark Database [GlobalBenchmarkDatabase], and observed that no system without extended-resolution reasoning could solve any benchmark in a time limit of 2 hours. We used a much larger timeout than with the two previous families because for this family we are not able to construct increasingly larger benchmarks and see how the solvers scale.
Table 1 reports on our experimental evaluation. In the first block we can see that the random and heuristic DIP choices do significantly improve upon our baseline solver. The second block row shows how the behavior of the heuristic configuration can be further improved by increasing the threshold of minimum DIP occurrences to 50. On the contrary, a smaller threshold of 5 significantly degrades the performance, whereas only learning the post-DIP clause slightly slows it down. Finally, the third block compares our best xMapleLCM configuration with state-of-the-art solvers. xMapleLCM and GlucosER are the only systems able to process any instance within the given time limit. Hence, the use of extended resolution is crucial in this set of benchmarks. Our comparison shows that GlucosER and xMapleLCM solve the same number of instances, the former being only slightly faster.
@@ 6.4. Performance Analysis on SAT Competition Benchmarks
In the left CDF plot of Figure 10 we can see that our baseline DIP-based xMapleLCM solver outperforms the CDCL-based MapleLCM that it is based on. When we analyzed the concrete benchmarks where our DIP-based solver outperforms MapleLCM and other CDCL-based solvers like Kissat, we observed that for all of them the percentage of decisions on extended variables by xMapleLCM is quite large (over 10% of the decisions), whereas for the remaining benchmarks it is very low (in 70% of the benchmarks less than 1% of decisions are on extended variables). Moreover, this happens no matter how many initial variables there are, even though in problems with few number of initial variables the introduction of a few extended variables could potentially quickly dominate the decision heuristic.
This is important since it shows that already existing decision heuristics like VSIDS or LRB somehow infer whether the newly introduced variables improve the behavior of the system. This is why we implemented, on top of our baseline DIP system, a procedure that computes the percentage of decisions on extended variables. If after a given number of conflicts it is still lower than 3%, it discontinues DIP learning and performs 1-UIP learning from that moment onwards. This is a very preliminary step in the direction of trying to automate the decision of whether to apply DIP reasoning or not, but the outcome, which is found on the left CDF plot of Figure 10, is very positive, showing that a solver disabling DIP reasoning (xMapleLCM-disabling in the plot) when the percentage of decisions on extended variables is low is still able to solve Tseitin, random $k$-xor and Interval benchmarks, but the overall performance is not significantly worsened on the remaining instances. Finally, the right CDF plot shows the impact of changing the type of DIP into the disabling solver: the heuristic and middle DIPs, the latter being used in the baseline configuration, are much better than random and closest ones.
@@
  • We have identified three families of benchmarks for which the dynamic addition of extension variables improves upon CDCL in a systematic manner. This is an improvement w.r.t. the results in the GlucosER paper [Audemard2010ARO], where a few initial benchmarks were identified but it was not proved that extended-resolution methods always outperform CDCL on increasingly more complex instances.
  • We have proved that, even though the way extended variables are identified is completely different in xMapleLCM and GlucosER, they both behave well on the same instances. On the other hand, our preliminary experiments revealed that randomly choosing arbitrary pairs of (non-DIP) literals to create definitions does not provide any speedup w.r.t. CDCL. Hence, both GlucosER and xMapleLCM are able to identify meaningful semantic properties about concrete literals that make them suitable for creating an extended variable definition. The data we collected show that GlucosER typically introduces around four times more extension variables than xMapleLCM and that the extended variables generated by GlucosER and xMapleLCM are neither disjoint nor totally overlap. Only about 25% of the definitions over original variables introduced in xMapleLCM are also introduced in GlucosER.
@@ 7. Conclusions and Future Work
We introduce a novel extended resolution clause learning (ERCL) algorithm that, when implemented on top of a CDCL solver, turns out to be beneficial for a variety of problems, in particular Tseitin, random $k$-xor and interval matching formulas. We view this as a step towards more effective methods for incorporating extended resolution to form more powerful proof systems. We show that the only previously existing attempt to incorporate ER into CDCL performs reasonably well on the same instances. Hence, we have indentified classes of formulas for which extended resolution can be automated in at least two different ways and still outperform CDCL. Considering the different nature of the two methods, this deserves further study. Further, we also introduce a new heuristic that allows our ERCL solver xMapleSAT to perform similarly to the baseline CDCL solver MapleLCM, thus being able to get the best of both worlds, i.e., the benefit of ERCL, without sacrificing performance of the CDCL solver on, say, SAT Competition Main Track 2023 instances.
As future work, we plan to investigate a variety of machine learning based heuristics (e.g., branching) specialized for the ERCL method. Also, the use of $k$-IPs with $k>2$ is part of our next steps. Finally, we intend to evaluate the impact of our technique on benchmarks from approximate hash-based counting, for which XOR reasoning is crucial. On the theoretical side, challenging questions like determining whether DIP- or $k$-IP based ER simulates unrestricted ER are going to be central to our research efforts.
Acknowledgements. We thank the anonymous referees for helpful comments and corrections.